OpenStudy (anonymous):

A room appearing to be rectangular has one side of length 408 cm, a second wall of length 306 cm, and one diagonal across the floor of length 525 cm.Is the room really rectangular? Explain

OpenStudy (amistre64):

I assume those are the measurements for a right triangle?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

408^2 + 306^2 = 525^2 .... see if this is a true statement :)

OpenStudy (amistre64):

260100 = 275625 not equal, not square, not rectangle

OpenStudy (amistre64):

I could tell by the last digits.. 8^2 + 6^2 = 5^2 64 + 36 = 25 100 = 25... the last digits 0 and 5 are not the same, so it wont fly no matter how large the numbers are...